High Sensitivity Detection of Trichloroethylene Using Frequency-Modulated Tunable Diode Lasers

\underline{Chi Kong Ng}, Khosrow Namjou, Edward A. Whittaker, Department of Physics and Engineering Physics, Stevens Institute of Technology, Hoboken, NJ 07030.

There is strong interest in the remote sensing community for high sensitivity optical sensors capable of distinguishing between various volatile organic compounds. In an effort to meet this need we have applied frequencymodulation technique absorption spectroscopy using tunable lead salt diode lasers to detect the compound trichloroethylene. Detection was accomplished using the relatively strong band at 1250cm^{−1}. While detection of a low pressure vapor was relatively straightforward, line blending in atmospheric pressure broadened samples causes much reduced sensitivity. We discuss possible approaches to solving this problem as well some specific details of our frequency modulation techniques which help to optimize achievable sensitivity.